Exercise 10-5 Working Backwards from Labor Variances [LO10-2]

The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

Standard Hours Standard Rate Standard Cost
Motor tune-up 1.90 $6.60 $12.54

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 200 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

Labor rate variance $ 954 F
Labor spending variance $ 36 U

Required:

1. Determine the number of actual labor-hours spent on tune-ups during the week.

2. Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.)
Answer #1
Labor efficiency variance = Labor spending variance+Labor rate variance
Labor efficiency variance = 36+954= $990 U
1
Labor efficiency variance = SR*(AH-SH)
990 = 6.60*(AH-200*1.90)
150 = AH-380
AH = 380+150 = 530
Actual labor-hours = 530
2
Labor rate variance = Actual hours*(SR-AR)
954 = 530*(6.60-AR)
1.80 = 6.60-AR
AR = 6.60-1.80 = 4.80
Actual hourly rate = $4.80